Can someone help me solve this quadtratic problem:
2x^2+7x+5

Found 2 solutions by shauna broeske, stanbon:
Answer by shauna broeske(4)
Answer by stanbon(75887)

solve this quadratic problem:
2x^2+7x+5=0
Factor:
2x^2+2x+5x+5 = 0
2x(x+1)+5(x+1) = 0
(x+1)(2x+5) = 0
x = -1 or x = -5/2
===============
Cheers,
Stan H>